OpenStudy (anonymous):

a plane flies 465 miles with the wind 315 miles against the wind in the same length of time, if the speed of the wind is 25mph, find the speed of the plane still in air

OpenStudy (anonymous):

r t = d, t = d/r \[\frac{465}{r+25}=\frac{315}{r-25}\]r=130mph

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Your Welcome.